using multiple steps solve following:
(6a-4)= 8
______
7

Answer by checkley77(12844)

(6a-4)= 8
______ CROSS MULTIPLY
7
6a-4=8*7
6a-4=56
6a=56+4
6a=60
a=60/6
a=10 ans.
Proof:
(6*10-4)/7=8
(60-4)/7=8
56=56